OpenStudy (anonymous):

A solid homogeneous cylinder of radius 'a' rolls without slipping inside a stationary hollow cylinder of large radius R. Write the Lagrangian and show that the motion of solid cylinder is simple harmonic.

OpenStudy (vincent-lyon.fr):

It is simple harmonic only at small angles.

OpenStudy (anonymous):

Oh.......... that i know!!!!!!!!

OpenStudy (vincent-lyon.fr):

Have you tried to write the kinetic energy and the potential energy of the solid? Use θ and dθ/dt to express them where θ is angle between vertical and line joining centre of big cylinder with centre of small cylinder. 

OpenStudy (anonymous):

\[K.E.= (1/2) m R^2\theta^.2 + (1/2) (I R^2 \theta^.2) /a^2\] Right??????

OpenStudy (vincent-lyon.fr):

Be careful that \(\omega\) of the cylinder is not \(\dot \theta\) and that radius of path of C is R-a.

OpenStudy (anonymous):

i got the answer \[\omega^2 = 2g/3(R-a)\] thanks. ur picture was really very much helpful. i was confused. but it's clear now

OpenStudy (vincent-lyon.fr):

What does this formula represent? \(\omega\) must depend on \(\dot \theta\) since the small cylinder is rolling inside the big one.

OpenStudy (vincent-lyon.fr):

The relation is: \(a\;\omega=-(R-a)\;\dot\theta\)

OpenStudy (anonymous):

yes i did like same and after that i got the final answer 'w'.

OpenStudy (vincent-lyon.fr):

I did not realise you meant \(\omega_o=2\pi/T_o\), pulsation of the oscillations.